OpenStudy (anonymous):

. In any circle, it takes_____ diameters to equal the circumference. A. about 3 2/7 B. exactly 3 1/7 C. about 3 1/14 D. about 3 1/7

OpenStudy (lgbasallote):

hint: \[\text{circumference} = \pi d\] now from your previous question what is the estimate of \(\pi\)

OpenStudy (anonymous):

pi = 22/7 = 3 1/7

OpenStudy (anonymous):

its pi and the value of pi is about 3 1/7